At Indianapolis Motor Speedway, one lap is 2.5 miles in length. The average speed of an Indy race car is 190 mph.
a.) Find the length of one lap in yards.
b.) How many seconds would it take to complete one lap?

Answer by RAY100(1637)

a) 2.5 miles *( 1760 yds/1 mi) =4400 yards
.
b),,,v=d/t,,,,or t= d/v= 2.5 m/ 190 MPH = .013 hr = .013hr(60min/1hr) = .789 min
.
or .789 min (60 sec / min) = 47.37 sec
.